Combined Communications is a new firm in a rapidly growing industry. The company is planning on increasing its annual dividend by 21 percent a year for the next 4 years and then decreasing the growth rate to 5 percent per year. The company just paid its annual dividend in the amount of $1.40 per share. What is the current value of one share of this stock if the required rate of return is 8.25 percent?
Answer #1

D1=(1.4*1.21)=1.694

D2=(1.694*1.21)=2.04974

D3=(2.04974*1.21)=2.4801854

D4=(2.4801854*1.21)=3.001024334

Value after year 4=(D4*Growth rate)/(Required rate-Growth rate)

=(3.001024334*1.05)/(0.0825-0.05)

=96.95617079

Hence current value=Future dividend and value*Present value of discounting factor(rate%,time period)

=1.694/1.0825+2.04974/1.0825^2+2.4801854/1.0825^3+3.001024334/1.0825^4+96.95617079/1.0825^4

=78.06(Approx).